OpenStudy (anonymous):

The point-slope formula of a line that passes through the point (8,5) and has a slope of ¼ is y = 1/4x + 3. A. True B. False

OpenStudy (anonymous):

solve it :) :\[y=mx+b\]\[5={1\over4}(8)+b\]\[5={8\over4}+b\]\[5=2+b\]\[3=b\]\[\large{y={1\over4}+3}\] \[\large\text{So in this case, it is 'True' :)}\]

OpenStudy (anonymous):

Ohhh

OpenStudy (anonymous):

Youre welcome